Solve the problem.
If $1250 is invested at a rate of 8% compounded monthly, what is the balance after 10 years?
[A = P(1 + i)n]

Answer by rfer(12731)

A=P(1+r/n)^n*t
A=1250(1+.08/12)^12*10
A=1250(1.00667)^120
A=1250*2.22
A=$2775